OpenStudy (anonymous):

Solve this inequality. -9x - 3 > 51 underscore >

OpenStudy (lgbasallote):

do you mean \[-9x - 3 \ge 51\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i mean x<=6

OpenStudy (anonymous):

how does this question different from other

OpenStudy (lgbasallote):

add 3 to both sides first \[-9x \ge 54\] divide both sides by -9 you get \[x \le -6\]

OpenStudy (anonymous):

oh.